OpenStudy (anonymous):

Please Helppp!!! Consider the sequence is 2,6,18,54 Let n = the term number in the sequence. Let A(n) = the vaule of the nth term of the sequence. What is the common ratio? i think its x's 3 Posting the table!

OpenStudy (whpalmer4):

6/2 = 3 18/6 = 3 54/18 = 3 3 would seem to be a good choice for the common ratio :-)